# Evaluating radiation impact on transmon qubits in above and underground facilities

### v1

@@ III Pulse Reconstruction
For the event detection, we developed an approach based on the probability of observing a certain stream of and . Initially, we performed a simultaneous Gaussian fit to the distribution of I/Q of , , and to correct for time instabilities and removed traces in which the qubit turned out to be in the second excited state . The clouds corresponding to and are then rotated in order to be aligned along one quadrature and projected onto the other. Further information can be found in the Supplementary Section D.
We then performed a Gaussian fit to determine the point at which and were maximally separated, enabling us to convert the measured I/Q values to a sequence of 0 (for ) and 1 (for ). This stream of zeros and ones was further processed to construct pulses for event detection. We started with the assumption that, for each measurement, the probability of the qubit remaining in state is equal to . Since, our readouts took several micro-seconds, we added that to the waiting period to accurately compute . Moreover, we found that the qubit relaxations varied by 25-30 over the data acquisition period (see Fig. 1(b)), consistent with the typical behavior of transmon devices [43]. Therefore, was computed using points (corresponding to approximately one minute worth of data) and used to estimate the probability for the qubit to be in state .
We began by picking the first 50 points of the trace and computing the probability for observing the measured counts of zeros and ones according to the Binomial probability distribution. Subsequently, we moved to the next point of the trace and repeated the calculation. Given that the probability for the qubit to be in state was approximately 85–90%, the probability of observing a large number of zeros in a subsequent series of measurements was very small. Finally, we computed the absolute logarithm () of this probability, simply to avoid working with very small numbers. If the probability of observing this series of events was smaller than 1 (), the software trigger was fired, and we saved a pulse with a length of 450 points, 100 of which were acquired before the pulse to monitor the status of the qubit before the interaction. Typical triggered pulses are shown in Fig. 4.
We then computed several quantities to characterize each pulse, including the value of the maximum, the length (defined as the number of points above threshold), the position of the maximum, as well as the value and root-mean-square (RMS) of the baseline before the pulse. These parameters were used to discard non-physical pulses (see orange trace in Fig. 4), such as those occurring on the tail of a preceding pulse or with a maximum too far from the trigger position. Consistency in data selection was ensured by applying the same criteria across all runs.
@@ IV Results and Discussion
The histogram of Max of the selected pulses is shown in Fig. 5. The single qubit measured in the fully shielded facility at LNGS features a rate of events/sec (solid gray trace). This is obtained by using the measurement time of 28 min, which is selected to filter periods with some unwanted excitation of the state. This shorter time causes larger statistical uncertainty. We repeated the analysis including longer intervals (in which state was sometimes excited) in order to increase the statistics and obtained events/sec (hatched gray trace). This factor two discrepancy can be considered as the systematic uncertainty on the result. The observed background event rate is about 2.5 to 5 times higher than the predicted rate from simulations, which is events/sec. This discrepancy suggests the presence of other sources of decoherence that resemble quasiparticle burst events but cannot be ascribed to radioactive sources.
Figure 5 clearly demonstrates that exposing the chip to a radioactive source results in an excess of events. The rate of these events increases linearly with the intensity of the source itself. Next, we compare the triggered and simulated rates for the various Thorium sources in Fig. 6.
@@
The slope parameter p1 of the linear fit in Fig. 6 shows that setting the threshold at 1 probability allows us to reach a detection efficiency of . This result indicates that qubits are sensitive to laboratory gamma radioactivity, despite the fact that gamma rays deposit a lower amount of energy compared to cosmic rays [22, 33]. This is not a significant concern for the development of quantum processors, as minimal passive shielding around the facility can largely mitigate this radioactive source [39, 50, 51]. Conversely, the absence of a high energy threshold for gamma detection presents interesting opportunities for using qubits as particle detectors [52]. While the primary objective of this work was not to operate a qubit as a particle detector, it is evident that optimizing the threshold or trigger algorithm, along with employing more refined pulse reconstruction algorithms (such as matched filter algorithms), could significantly improve detection efficiency.
However, from the linear fit the intercept p0 results in () events/second. This parameter can be interpreted as the expected rate in the absence of radioactive interactions (if the qubit response to the source is fully linear) and exceeds the one measured in the “background” run. Future work will investigate this discrepancy further.
We deepened the study of the effect of radioactivity by investigating a possible activation of two-level-systems (TLS) when deploying the gamma source. As pointed out in Ref. 23, the ionizing effect of radiation on the chip and the consecutive charge recombination could locally affect the distribution of charges and the electric field that they produce, to which TLSs are sensitive. As a result, the TLSs’ frequencies could shift, getting into resonance or out of resonance with the qubit. This would cause a fluctuation of the qubit T1 lasting longer than the recombination time of the quasiparticles produced by an impact.
We analyzed the data obtained from the measurements done at LNGS when exposing the qubit to the 161 kBq Thorium source. We selected pulses with very high maxima (), so to discard possible noise events. For each pulse we calculated the average values of approximately 10 ms before the trigger and tens of milliseconds after it, when the quasiparticle poisoning was already over. A change in the qubit T1 due to TLS dynamics would result in an average value of after the pulse different from what obtained before the trigger.
We made the histogram of for radioactivity impacts and compared it with what obtained by selecting noise events instead (Fig. 7). We didn’t observe significant differences between the two histograms. To double check this result we exposed our qubit to a Cesium source which activity was an order of magnitude higher than the Th source one. Also in that case, no significant differences between radioactivity events and noise events were observed.
This analysis showed that the fluctuations in the qubit T1 correlated to radioactivity impacts were of the same magnitude of what caused by other sources of noise. Given also the high rate of noise events that we observed, we concluded that ionizing radiation does not play a major role in undermining the T1 stability of the qubit we studied.
Finally, we investigated whether radioactive impacts could excite the qubit from to state. To this aim, we performed the same study described above, but with a reversed measurement protocol: initializing the qubit in the state instead of , and monitoring the qubit exposed to a strong gamma source. Interestingly, exposing the qubit to gamma radiation did not result in an excess of events as demonstrated in Fig. 10. This observation is consistent with what was previously reported in Ref. 20, confirming that quasiparticles generated due to radiation do not deliver energy to transmon qubits.
Finally, we conducted a similar fast decay protocol study using multiple “twin-chips” operated at the above ground facilities at SQMS Fermilab. Here qubits are fully unshielded from cosmic and gamma radiation. The results showed consistency with those obtained at LNGS, as we obtained qualitatively similar distributions in all the pulse parameters defined for our study. The distribution of Max in particular, is reported in Figure 8. The histograms do not reveal any distinct behavior of the chip in the measurements conducted deep underground at LNGS compared to those above-ground at FNAL. The protocol applied to three qubits at FNAL under ambient radiation conditions showed varied responses: similar (Q4), reduced (Q2), and increased (Q3) rates when compared to the response of the maximally shielded qubit at LNGS.
These variations observed in the rates for different qubits measured at FNAL compared to the LNGS background one indicate that modern transmons are dominated by other sources of noise and radiation does not play a major role in the drastic drop in T1 at short time scales. This fact has also been corroborated by another recent study [41] where cosmic rays are not found to be a strong contributor to the most frequent correlated error (four qubits).
Even though we did not experimentally observe a measurable effect of environmental radiation induced errors, we can utilize the simulated rates to give a cap estimate for the relevant timescales at which radiation would start to affect quantum computation.
@@
The probability to have at least one event in a time window T is given by , where is the rate of impacts. Using the simulated rates, for environmental radioactivity we find that in a non-shielded laboratory setting a quantum computation would be affected by radioactivity less than 0.1% (1%) times if the computation is completed within approx (). Note that the computation may include multiple runs of an algorithm and not necessarily require to be a single run lasting for T. At a fully shielded environment like LNGS, due to much lower expected rate, we have instead (1%) if (2.5 s). The available computation period increases if higher error probability is tolerated. Figure 9 shows available computational time as a function of radiation impact rates around values estimated for the two cases. The contour lines show various constant error rates and the horizontal green shaded area ( ms) represents the T1 times of contemporary transmon qubits. In this regime, radiation causes error with a minuscule . In other words, when a calculation is repeated times, one result will be corrupted due to radiation. Most of the modern algorithms running on quantum hardware like variational quantum eigensolver or quantum approximate optimization take less than a milli-second for a single iteration and require a large number of repetitions (). Therefore, we can safely conclude that radiation has negligible effect on modern transmon-based quantum computing platforms even if error-correction fails due to an impact.
@@ V Conclusions
We have performed a comparative study of how radiation may affect transmon qubits hosted at two laboratories with very distinct radiation environments. The regular T1 experiments performed on the same qubit showed no difference in qubit lifetimes of the order of 100 microseconds. The experiments with controlled Thorium sources conducted at the deep underground Gran Sasso laboratory demonstrate the possibility of achieving a good detection efficiency, opening the possibility of using transmons as a particle detector. However, the background rate observed with no active sources at LNGS show higher event rates compared to the simulated values, indicating that other noise sources produce signatures similar to those expected or conjectured to be coming from radiation.
The same protocol when applied to multiple similar qubits in Fermilab (above-ground and with no lead shield) with ambient radiation, shows both increased and reduced rates compared to the response of the maximally shielded qubit at LNGS. These results suggest that these radiation-like events lasting milliseconds are generated by different sources, and that radiation is not a major source of errors in modern transmon qubits. Based on the experimental and simulated results we find that quantum computation will be affected at a maximum of a percent level due to radiation inflicted single-qubit errors when computational time reaches more than 170 ms at fully unshielded facilities and 2.5 s at LNGS. We found no correlation of radiation events to TLS scrambling or qubit excitation.
Future experiments will aim at in-depth understanding of the various error sources which are currently dominating over the radiation induced effects. This will involve further investigations based on fast decay protocols testing different qubit geometries, materials and synchronized measurements across same and different chips. SQMS is focused on enhancing the T1 of transmon qubits and so future measurements will investigate the behavior of qubits with a wide distribution of coherence performance ranging from tens of microseconds to approaching milliseconds. Advancing this knowledge will be critical for the use of superconducting qubits for quantum computation and sensing.
@@ Acknowledgements.
The authors thank the Director and technical staff of the Laboratori Nazionali del Gran Sasso. We are also grateful to the LNGS Computing and Network Service for computing resources and support on U-LITE cluster at LNGS. We are grateful to A. Girardi for the controller of the cryogenic switch, M. Guetti for the help with the whole cryogenic facility (including its continuous upgrades), and M. Iannone for the assembly of the device at LNGS. We are thankful to Theodore C. III and Grzegorz Tatkowski for providing support on experimental preparations and cryogenic operations at FNAL. T.R. thanks Xinyuan You for useful discussions. The work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems (SQMS) Center under the contract No. DE-AC02-07CH11359, by the Italian Ministry of Foreign Affairs and International Cooperation, grant number US23GR09, by the Italian Ministry of Research under the PRIN Contract No. 2020h5l338 (“Thin films and radioactivity mitigation to enhance superconducting quantum processors and low temperature particle detectors”) and by the Italian Ministry of Research under the PRIN Contract No. 2022BP4H73 (Chasing Light Dark Matter with Quantum Technologies”).References
@@
- Fowler et al. [2024] J. W. Fowler, P. Szypryt, R. Bunker, E. R. Edwards, I. F. Florang, J. Gao, A. Giachero, S. F. Hoogerheide, B. Loer, H. P. Mumm, N. Nakamura, G. C. O’Neil, J. L. Orrell, E. M. Scott, J. Stevens, D. S. Swetz, B. A. VanDevender, M. Vissers, and J. N. Ullom, Spectroscopic measurements and models of energy deposition in the substrate of quantum circuits by natural ionizing radiation (2024), arXiv:2404.10866 .
- Loer et al. [2024] B. Loer, P. M. Harrington, B. Archambault, E. Fuller, B. Pierson, I. Arnquist, K. Harouaka, T. D. Schlieder, D. K. Kim, A. J. Melville, B. M. Niedzielski, J. K. Yoder, K. Serniak, W. D. Oliver, J. L. Orrell, R. Bunker, B. A. VanDevender, and M. Warner, Abatement of ionizing radiation for superconducting quantum devices (2024), arXiv:2403.01032 .
- Linehan et al. [2024] R. Linehan, I. Hernandez, D. J. Temples, S. Q. Dang, D. Baxter, L. Hsu, E. Figueroa-Feliciano, R. Khatiwada, K. Anyang, D. Bowring, G. Bratrud, G. Cancelo, A. Chou, R. Gualtieri, K. Stifter, and S. Sussman, Estimating the energy threshold of phonon-mediated superconducting qubit detectors operated in an energy-relaxation sensing scheme (2024), arXiv:2404.04423 .

### v2

@@ III Event Selection
Several considerations are applied to the analysis for the selection of events. Initially, we perform a simultaneous 2D Gaussian fit to the distribution of I/Q values for the states , , and . To correct for time instabilities, we remove traces where the qubit is found in the second excited state . The clouds corresponding to and are then rotated to align along one quadrature and projected onto the other. Further information can be found in the Supplementary Section D.
Once the I/Q clouds are projected onto one quadrature, a threshold is chosen in order to convert the measured I/Q values into a sequence of 0s (for ) and 1s (for ). Since the two clouds are not completely separated, every value of the threshold will lead to some errors in the state identification. In our analysis, we calculate the fraction of measurements for each state falling on the wrong side of the threshold (misidentification) and choose the threshold value that minimizes the sum of these two fractions (total error). This is done by fitting the clouds with 1D Gaussian functions and integrating the tails falling on the wrong side of the threshold, varying the candidate threshold value until the minimum is found. The resulting threshold values allow for the correct identification of the qubit state in 90% to 95% of measurements, with fluctuations of a few percentage points depending on the specific qubit and dataset considered.
The sequence of 0s and 1s is then organized into traces of points. We expect that radiation-induced events last at least one millisecond, which, given the sampling periods reported in Table 3, corresponds to approximately 12–35 zeros. According to the same table, we anticipate a baseline of 8–15% of zeros arising from spontaneous qubit decay. Therefore, we require the detection of at least four consecutive zeros to trigger an event. This criterion ensures reasonable signal efficiency while significantly reducing false triggers caused by qubit spontaneous decay and other types of noise. Indeed, the number of times we measure the qubit in state can be due to different phenomena, which can be summarized as follows:
1.
Qubit spontaneous decay: The probability of finding the qubit in state can be roughly estimated by , where is an “effective” relaxation time computed using points (corresponding to approximately one minute of data) to monitor changes over time, and is the waiting period (see Table 3) between the state preparation in and end of readout. We then use the average of the values obtained from each data segment to estimate the probability of the qubit being in state for each dataset. Further details can be found in Supplementary Section E.
2.
Ionizing radiation: We use Geant4-based Monte Carlo simulations to predict the number of impacts on the qubit chip.
3.
Other sources of noise: These involve electronic noise, luminescence or relaxation of materials at low temperature [51] (known sources of quasiparticle poisoning include the holder or the interface between the holder and the chip [52, 53, 54] and relaxations of metal films [55, 53]), vibrations [56], etc.
Datasets | TS | r | |||
[s] | [s] | [%] | [s] | [events/sec] | |
Shielded chip | 14 | 118.6 | 11.1 | 73.7 | 2.3 0.4 |
44 kBq Source | 11 | 76.3 | 13.4 | 67.7 | 5.4 1.1 |
76 kBq Source | 11 | 79.9 | 12.9 | 67.7 | 4.6 0.9 |
125 kBq Source | 11 | 68.9 | 14.8 | 67.7 | 7.2 1.5 |
161 kBq Source | 11 | 74.1 | 13.8 | 67.7 | 6.1 1.2 |
FNAL Q2 | 14 | 162.7 | 8.2 | 88.0 | 0.5 0.1 |
FNAL Q3 | 9 | 63.8 | 13.2 | 84.3 | 4.1 0.8 |
FNAL Q4 | 13 | 165.5 | 7.6 | 31.7 | 1.1 0.2 |
Once the trigger is fired, we acquire a 160-point window divided as follows:
A 60-point long () pre-trigger region to estimate the average number of zeros (noise) before the events;
A 40-point long () dead region to avoid signal leakage into the pre-trigger region;
A 60-point long () signal region. Specifically, we acquire 10 points before the trigger and 50 after the trigger.
The choice of the window length was based on the signal features. Given a sampling period of 0.03–0.09 milliseconds, and considering that an event lasts a few milliseconds (see Fig. 1), we expect the event to be fully contained within a 60-point window. To prevent re-triggering, we apply a dead-time of 51 points after the trigger.
As shown in Table 3, we expect from 0.5 up to 7 events per second due to qubit spontaneous decay alone, depending on the dataset. These rates are about one to two orders of magnitude higher than those expected from ambient radiation (0.004 and 0.057 events/sec at LNGS and FNAL, respectively) or gamma ray sources (0.12 to 0.43 events/sec).
To identify events caused by ionizing radiation, we employ a systematic approach aimed at minimizing contributions from qubit spontaneous decay. For each dataset, we compute the probability of observing more than a certain number of zeros in a sequence of 60 measurements, corresponding to the length of the signal window. This probability is derived from the cumulative distribution function (CDF) of the Binomial distribution calculated over 60 points. Based on these probabilities, we set thresholds on the minimum number of zeros required within the signal region to significantly reduce the expected rate of qubit spontaneous decay, targeting values approximately an order of magnitude lower than those expected from ionizing radiation.
Additionally, we establish thresholds on the maximum number of zeros allowed in the pre-trigger region. In this region, the number of zeros should be consistent with those predicted by the qubit decay. A significantly larger number of zeros would indicate transient noise that could induce false triggers and has to be discarded from the analysis. Therefore, we calculate the Binomial probability of observing a specific number of zeros in the pre-trigger region and apply a cut for probabilities below 1%. Detailed thresholds used for each dataset are provided in Supplementary Section F.
@@ IV Results and Discussion
The distributions of the number of zeros in the signal region for the triggered events are depicted in Fig. 4. These distributions correspond to background runs conducted at the fully shielded facility at LNGS without any Thorium source and at FNAL, where qubits were entirely exposed to cosmic rays and gamma radiation in the laboratory environment. It can be observed that the histograms of Q1 and Q3, which have similar values, also exhibit similar shapes. This observation holds true for Q2 and Q4 as well.
As described in the previous section, to identify radiation-induced events, we reduce qubit spontaneous decay events to negligible rates by setting a threshold on the minimum number of zeros in the signal region. For example, for the LNGS–Q1 and FNAL–Q3 datasets, which feature a similar , we required at least 23 zeros (non-consecutive) in order to have a negligible rate from qubit spontaneous decay. In contrast, for FNAL–Q2 and FNAL–Q4, we required at least 18 zeros due to the lower . We emphasize that such a stringent requirement discards the vast majority of the events (see Fig. 4).
Additionally, a maximum number of zeros in the pre-trigger region is set to mitigate noisy periods. This number ranges from a minimum of 10, for datasets with lower to a maximum of 15 for datasets with higher . The cuts on the number of zeros in the pre-trigger and in the signal regions for all the datasets are reported in Table S4 of the Supplement. Table 4 summarizes the results obtained from the background runs conducted at LNGS and FNAL. The three qubits operated on the same chip at FNAL show rates ranging from 3 to 2510-3 events/sec, in contrast to what we would expect if the selected events were dominated by radioactivity. In this case, indeed, the rates should have been similar. Moreover, the rate measured in the fully shielded facility at LNGS is comparable with the one measured at FNAL. If the obtained values were ascribed to radioactivity, the LNGS rate should have been an order of magnitude smaller than the FNAL one. These results suggest that modern transmons are primarily affected by sources of noise other than radiation, contributing significantly to the reduction in T1 at short timescales. This fact has also been corroborated by another recent study [38], where cosmic rays were not identified as a major contributor to the most frequent correlated errors (among four qubits).
Finally, we observe that the measured rates are lower than those expected from simulations of ambient radiation. This discrepancy suggests that the event selection substantially reduces the efficiency of radiation-induced events.
Datasets | Measured Rate | Radioactivity Contribution |
[events/sec] | [events/sec] | |
Shielded chip | (2.9 1.3) | (4.0 0.6) |
FNAL Q2 | (3.2 0.7) | |
FNAL Q3 | (4.7 0.9) | (57 3) |
FNAL Q4 | (25 4) |
To better investigate the detection efficiency for radiation-induced events, at LNGS, we exposed the chip to Thorium radioactive sources with increasing activity levels. Several commercially available Thorium rods with known activity were inserted inside the copper shield and outside the outer vacuum chamber of the fridge for this study. Data were analyzed using the same protocol used for background runs and reported in Fig. 5. It demonstrates that exposing the chip to a controlled radioactive source results in an excess of events with high number of zeros in the signal region compared to the “shielded” configuration. The rate of these events increases linearly with the intensity of the source itself as visible in Fig. 6.
Next, we compare the triggered and simulated rates for the various Thorium sources. The slope parameter p1 of the linear fit in Fig. 6 shows that the aggressive data selection leads to a detection efficiency of , where the uncertainty is statistic only. The systematic uncertainty is much larger, as the signal efficiency depends on the cuts used for data selection, that vary across the datasets because of the differing and noise conditions. Therefore, this result should be considered qualitative. Correcting the predicted rate at LNGS (410-3 event/sec) by this efficiency results in an expected rate from radioactivity of 0.3210-3 event/sec, an order of magnitude smaller than the measured one. This further supports the hypothesis that the measured rates are not dominated by radiation-induced events.
@@
Despite the low detection efficiency, our findings indicate that qubits are sensitive to laboratory gamma radioactivity, even though gamma rays deposit less energy compared to cosmic rays [22, 34]. While the primary objective of this work was not to operate a qubit as a particle detector, it is evident that optimizing the thresholds or trigger algorithms, along with employing more refined pulse reconstruction techniques (such as matched filter algorithms), could significantly improve detection efficiency.
@@
Although we did not experimentally observe a measurable effect of environmental radiation-induced errors, we can utilize the simulated rates to provide an upper estimate for the relevant timescales at which radiation would start to affect quantum computation through information loss occurring on individual qubits. The probability to have at least one event in a time window T is given by , where is the rate of impacts. Using the simulated rates, for environmental radioactivity we find that in a non-shielded laboratory setting a quantum computation would be affected by radioactivity less than 0.1% (1%) times if the computation is completed within approx (). Note that the computation may include multiple runs of an algorithm and not necessarily require to be a single run lasting for T. At a fully shielded environment like LNGS, due to a much lower expected rate, we have instead (1%) if (2.5 s). The available computation period increases if a higher error probability is tolerated. Figure 7 shows available computational time as a function of radiation impact rates around values estimated for the two cases. The contour lines show various constant error rates and the horizontal green shaded area ( ms) represents the T1 times of contemporary transmon qubits. In this regime, radiation causes error with a minuscule . In other words, when a calculation is repeated times, one result will be corrupted due to radiation. Most of the modern algorithms running on quantum hardware like variational quantum eigensolver or quantum approximate optimization take less than a millisecond for a single iteration and require a large number of repetitions (). Therefore, we can safely conclude that radiation has a negligible effect on modern transmon-based quantum computing platforms even if error correction fails due to an impact.
@@ V Conclusions
We have conducted a comparative study to assess the impact of radiation on transmon qubits at two laboratories with very different radiation environments. Regular T1 experiments performed on the same qubit show no significant difference in qubit lifetimes of the order of 80 microseconds.
In order to study the time dynamics at short timescales, we employed a fast decay detection protocol utilizing active reset. Multiple similar qubits measured with this technique at Fermilab (above ground and without lead or copper shieldings) showed rates comparable to those from the maximally shielded qubit at the deep-underground Gran Sasso Laboratory. These findings suggest that the radiation-like events observed, which last for milliseconds, originate from different sources, and indicate that radiation is not a major source of errors in modern transmon qubits.
Experiments with controlled Thorium sources at LNGS demonstrate the potential for detecting particle impacts, opening the possibility of using transmons as particle detectors. However, the observed rates are lower than those predicted by simulations, indicating the need for more refined trigger algorithms and/or analysis strategies.
Based on both experimental and simulated results, we estimate that radiation-induced single-qubit errors will affect quantum computation by at most a few percent, with computational times up to 170 ms at fully unshielded facilities and 2.5 s at LNGS.
Future experiments will focus on a more detailed understanding of the various error sources currently overshadowing radiation-induced effects. This will involve further investigations using fast decay detection protocols, exploring different qubit geometries, materials, and synchronized measurements across the same and different chips. The SQMS Center aims to enhance the T1 of transmon qubits, and future studies will examine the performance of qubits with a broad range of coherence times, from tens of microseconds to approaching milliseconds. Advancing this knowledge is crucial for the effective use of superconducting qubits in quantum computation and sensing.
@@ Acknowledgements.
The authors thank the Director and technical staff of the Laboratori Nazionali del Gran Sasso. We are also grateful to the LNGS Computing and Network Service for computing resources and support on U-LITE cluster at LNGS. We thank S. Nisi for the ICP-MS measurements of the PCB, C. Tomei and G. D’Imperio for the maintenance of the simulation software, G. Catelani for the useful discussion. We are grateful to A. Girardi for the controller of the cryogenic switch, M. Guetti for the help with the whole cryogenic facility (including its continuous upgrades), and M. Iannone for the assembly of the device at LNGS. We are thankful to Theodore C. III and Grzegorz Tatkowski for providing support on experimental preparations and cryogenic operations at FNAL. The work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems (SQMS) Center under the contract No. DE-AC02-07CH11359, by the Italian Ministry of Foreign Affairs and International Cooperation, grant number US23GR09, by the PNRR MUR project CN00000013-ICSC, by the Italian Ministry of Research under the PRIN Contract No. 2020h5l338 (“Thin films and radioactivity mitigation to enhance superconducting quantum processors and low temperature particle detectors”) and by the Italian Ministry of Research under the PRIN Contract No. 2022BP4H73 (Chasing Light Dark Matter with Quantum Technologies”).References
@@
- Bratrud et al. [2024] G. Bratrud, S. Lewis, K. Anyang, A. Cesaní, T. Dyson, H. Magoon, D. Sabhari, G. Spahn, G. Wagner, R. Gualtieri, et al., First measurement of correlated charge noise in superconducting qubits at an underground facility, arXiv preprint arXiv:2405.04642 (2024).
@@
- Adari et al. [2022] P. Adari et al., Excess workshop: Descriptions of rising low-energy spectra, SciPost Physics Proceedings 10.21468/scipostphysproc.9.001 (2022).
- Anthony-Petersen et al. [2022] R. Anthony-Petersen et al., A stress induced source of phonon bursts and quasiparticle poisoning (2022), arXiv:2208.02790 [physics.ins-det] .
- Angloher et al. [2023] G. Angloher et al., Latest observations on the low energy excess in cresst-iii, SciPost Physics Proceedings 10.21468/scipostphysproc.12.013 (2023).
- Armengaud et al. [2016] E. Armengaud et al., Constraints on low-mass wimps from the edelweiss-iii dark matter search, Journal of Cosmology and Astroparticle Physics 2016 (05), 019–019.
- Romani [2024] R. K. Romani, Aluminium relaxation as the source of excess low energy events in low threshold calorimeters (2024), arXiv:2406.15425 [physics.ins-det] .
- Kono et al. [2024] S. Kono, J. Pan, M. Chegnizadeh, X. Wang, A. Youssefi, M. Scigliuzzo, and T. J. Kippenberg, Mechanically induced correlated errors on superconducting qubits with relaxation times exceeding 0.4 ms, Nature Communications 15, 3950 (2024).
- Verkerke and Kirkby [2003] V. Verkerke and D. Kirkby, The roofit toolkit for data modeling (2003), arXiv:0306116 .